Voter turnout in the preceding 10 years can be modeled by the function N(t) = 12432 + 11t - t2 where N is the number of voters and t is measured in years. What was the rate of change in the number of voters for year 8?
Find the average rate of change for f(x) = 2x - x3 from x = 3 to x = 8. (2 decimal places)

Answer by stanbon(75887)

Voter turnout in the preceding 10 years can be modeled by the function N(t) = 12432 + 11t - t^2 where N is the number of voters and t is measured in years. What was the rate of change in the number of voters for year 8?
Find the average rate of change for f(x) = 2x - x^3 from x = 3 to x = 8. (2 decimal places)
----------
N'(t)=11-2t
N'(8)=11-2(8)=11-16=-5
---------------
Ave rate of change = [f(2)-f(3)]/[2-3]
=[(2*2-2^3)-(2*3-3^3)]/(-1)
=[-4+3]/(-1)
=1
Cheers,
Stan H.